# Holmes: Towards Effective and Harmless Model Ownership Verification to Personalized Large Vision Models via Decoupling Common Features

### v1

@@ V-D3 The Effectiveness of the Selection Strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$
Distillation 0.18 $\mathbf{10^{-12}}$ 0.01 $10^{-3}$ 0.11 $10^{-6}$
DFKD 0.13 $\mathbf{10^{-9}}$ 0.08 $10^{-3}$ 0.09 $10^{-5}$
Label-query 0.05 $\mathbf{10^{-4}}$ 0.01 0.01 0.03 0.01
Logit-query 0.14 $\mathbf{10^{-8}}$ 0.02 $10^{-3}$ 0.08 $10^{-3}$
Independent -0.03 0.20 $10^{-3}$ $10^{-3}$0.01 0.12
Our method selects $\gamma\%$ of samples with the lowest loss values to form $\mathcal{D}_{s}$ and drops $\lambda\%$ with the lowest loss values to form $\mathcal{D}_{f}$, as these samples indicate better learning by the victim model and a higher degree of overfitting. Consequently, they are more representative of dataset-specific features. To verify the effectiveness of this selection strategy, we conduct experiments using two alternative approaches: selecting the $\gamma\%$ ($\lambda\%$) of samples with the highest loss values and randomly selecting $\gamma\%$ ($\lambda\%$) of the samples.
As shown in Table XI-XII, selecting samples with the lowest loss values yields the best verification results in almost all cases for both $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$ selections. For $\mathcal{D}_{s}$, selecting samples with the highest loss values fails in cases of fine-tuning attack, distillation attack, and identifying the independent model. This failure is probably because the samples with the highest loss values are less effectively learned by the victim model compared to other samples, making them inadequate to accurately capture dataset-specific features, leading to ineffective verification. For the $\mathcal{D}_{f}$ selection, dropping the highest-loss samples causes failure in identifying the independent model. This is probably because dropping highest-loss samples (instead of dropping lowest-loss ones) makes the benign shadow model retain dataset-specific features too similar to those of the victim model. Consequently, the meta-classifier struggles to distinguish independent models that exhibit similar but distinct dataset-specific features, leading to misjudgment.
V-E Resistance to Potential Adaptive Attacks
In this section, we analyze the resistance of our method against potential adaptive attacks. Specifically, we consider overwriting attack, unlearning attack, and model pruning [63] as possible attack strategies, where the adversary is aware of our defense method and has access to $\mathcal{D}_{s}$ and $\mathcal{D}$.
V-E1 Resistance to Overwriting Attack
The adversary may attempt to overwrite the dataset-specific features inherited from the victim model by implanting alternative features through fine-tuning. Specifically, the adversary fine-tunes the victim model for 20 epochs on the TinyImages dataset (distinct from the victim’s CIFAR-10 training data) to assess the erasure of original dataset-specific features. As shown in Table XIII, our method remains effective against the overwriting attack. These results demonstrate that even when adversaries are aware of our defense method but lack specific implementation details, our method robustly resists adaptive fine-tuning designed to eliminate dataset-specific features.
V-E2 Resistance to Unlearning Attack
TABLE XIII: The resistance to adaptive attacks.
@@ V-E2 Resistance to Unlearning Attack
Distillation 0.18 $10^{-12}$ 0.22 $10^{-7}$ 0.14 $10^{-7}$ 0.10 $10^{-8}$
DFKD 0.13 $10^{-9}$ 0.04 $10^{-3}$ 0.01 0.01 0.09 0.01
Label-query 0.05 $10^{-4}$ 0.03 $10^{-4}$ 0.01 $10^{-3}$ 0.01 0.01
Logit-query 0.14 $10^{-8}$ 0.03 $10^{-4}$ 0.14 $10^{-4}$ 0.08 $10^{-3}$
Independent -0.03 0.20 -0.01 0.10 -0.02 0.19 -0.02 0.12
@@
In this scenario, the adversary attempts to make the victim model forget the dataset-specific features, thereby undermining the effectiveness of model ownership verification. Specifically, the adversary randomly reassigns the labels of all samples in $\mathcal{D}_{s}$ to incorrect labels, creating a modified dataset $\mathcal{D}^{\prime}_{s}$. The adversary then fine-tunes the victim model on $\mathcal{D}^{\prime}_{s}\cup\mathcal{D}_{b}$ for 20 epochs, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$. As shown in Table XIII, our method remains effective in defending against the unlearning attack, demonstrating its robustness.
V-E3 Resistance to Model Pruning
@@
Given the prevalent use of model pruning in adapting personalized LVMs, we evaluate its potential impact on our method’s verification reliability. In this experiment, we apply retraining-free pruning [63] to the victim model, following the original paper’s settings. As shown in Table XIII, our method still succeeds in all cases, demonstrating its robustness.
VI Potential Limitations and Future Directions
@@
Although our method demonstrates robust performance against various model stealing attacks (Section V), we acknowledge several limitations that warrant discussion.
@@
Indirect Representation of Dataset-Specific Features. Our method utilizes a backdoor attack to disrupt the dataset-specific features learned by the victim model. While our experiments show that this approach is effective, we acknowledge that other techniques could be explored for decoupling common features and disrupting dataset-specific ones.
@@
Computational Overhead. Our method involves fine-tuning a poisoned shadow model, a benign shadow model, and training an external meta-classifier, which introduces additional computational overhead. However, training shadow models for reference is a common practice in model watermarking and fingerprints to evaluate whether these techniques impact the model’s performance [17, 25, 18]. In our method, the fine-tuning cost of the shadow models is comparable to that of fine-tuning the victim model itself. For example, in the main experiments in Section V, fine-tuning a shadow model on the CIFAR-10 dataset takes 1 hour and 58 minutes, while fine-tuning on the ImageNet dataset takes 1 hour and 7 minutes, both using an NVIDIA A100 GPU. Besides, this process can be accelerated through parallel GPU execution. The training cost for the meta-classifier is relatively mild, as its structure is lightweight and requires only a limited number of training samples. For example, training it on CIFAR-10 and ImageNet takes approximately 5 minutes on an NVIDIA A100 GPU.
@@
Generalization to Diverse Data Domains. While our method demonstrates effectiveness on standard vision benchmarks (CIFAR-10 and ImageNet), its robustness on specialized domains ($e.g.$, medical imaging, low-resolution satellite imagery) remains unverified. The variations in data characteristics of these specialized domains, such as texture patterns or semantic granularity, might affect the decoupling of common and dataset-specific features, warranting further investigation.
VII Conclusion
@@
In this paper, we revisited the challenge of verifying ownership for personalized large vision models. We revealed that existing methods, which are typically designed for models trained from scratch, are either ineffective for fine-tuned models, introduce additional security risks, or are prone to misjudgment. To address these issues, we proposed a harmless model ownership verification method by decoupling common features. By creating a poisoned shadow model (retaining common features but disrupting dataset-specific ones via backdoor attacks) and a benign shadow model (learning distinct dataset-specific features), we distinguished the victim model’s dataset-specific features using output differences. A meta-classifier and hypothesis test further enhanced verification reliability. We evaluated our method on the CIFAR-10 and ImageNet datasets, and the results demonstrated its effectiveness in detecting various types of model stealing attacks simultaneously.

### v2

@@ 5.4.3 The Effectiveness of the Selection Strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$
Data Distill 0.18 $\mathbf{10^{-12}}$ 0.01 $10^{-3}$ 0.11 $10^{-6}$
Data-free Distill 0.13 $\mathbf{10^{-9}}$ 0.08 $10^{-3}$ 0.09 $10^{-5}$
Hard Distill 0.05 $\mathbf{10^{-4}}$ 0.01 0.01 0.03 0.01
Soft Distill 0.14 $\mathbf{10^{-8}}$ 0.02 $10^{-3}$ 0.08 $10^{-3}$
Independent -0.03 0.20 $10^{-3}$ $10^{-3}$ -0.01 0.12
Table 13: The resistance to potential adaptive attacks.
@@
[…]
Data-free Distill 0.13 $10^{-9}$ 0.04 $10^{-3}$ 0.01 0.01 0.09 0.01
Hard Distill 0.05 $10^{-4}$ 0.03 $^{-4} 0.01 1$$^{-3} 0.01$ 0.01
Soft Distill 0.14 $10^{-8}$ 0.03 $10^{-4}$ 0.14 $10^{-4}$ 0.08 $10^{-3}$
6 Discussion
6.1 Resistance to Potential Adaptive Attacks
In this section, we analyze the resistance of our method against potential adaptive attacks. Specifically, we consider overwriting attack, unlearning attack, and model pruning [kwon2022fast] as possible attack strategies, where the adversary is aware of our defense method and has access to $\mathcal{D}_{s}$ and $\mathcal{D}$.
6.1.1 Resistance to Overwriting Attack
The adversary may attempt to overwrite the dataset-specific features inherited from the victim model by implanting alternative features through fine-tuning. Specifically, the adversary fine-tunes the victim model for 20 epochs on the TinyImages dataset (distinct from the victim’s CIFAR-10 training data) to assess the erasure of original dataset-specific features. As shown in Table 13, our method remains effective against the overwriting attack. These results demonstrate that even when adversaries are aware of our defense method, our method robustly resists adaptive fine-tuning designed to eliminate dataset-specific features.
6.1.2 Resistance to Unlearning Attack
@@
In this scenario, the adversary attempts to make the victim model forget the dataset-specific features, thereby undermining the effectiveness of model ownership verification. Specifically, the adversary randomly reassigns the labels of all samples in $\mathcal{D}_{s}$ to incorrect labels, creating a modified dataset $\mathcal{D}^{\prime}_{s}$. The adversary then fine-tunes the victim model on $\mathcal{D}^{\prime}_{s}\cup\mathcal{D}_{b}$ for 5 epochs, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$. As shown in Table 13, our method remains effective in defending against the unlearning attack, demonstrating its robustness.
6.1.3 Resistance to Model Pruning
@@
Given the prevalent use of model pruning in adapting personalized LVMs, we evaluate its potential impact on our method’s verification reliability. In this experiment, we apply retraining-free pruning [kwon2022fast] to the victim model, following the original paper’s settings. As shown in Table 13, our method still succeeds in all cases, demonstrating its robustness against pruning.
6.2 Potential Task Adaptation
Arguably, our Holmes framework is inherently adaptable to more complex visual tasks beyond image classification (e.g., image captioning). In this section, we use image captioning as a representative complex generative task to discuss how Holmes can be extended accordingly.
6.2.1 Method Adaptation
When adapting to the image captioning task, the core three-stage pipeline of our Holmes (shadow model construction, meta-classifier training, hypothesis testing) remains unchanged, with key modifications limited to adjusting poisoned shadow model construction and projecting text outputs into the feature space.
Shadow Model Construction for Image Captioning. In general, shadow model construction follows the same principle in Section 4.1, involving a poisoned shadow model $P_{\text{cap}}$ (preserving common features while disrupting dataset-specific features) and a benign shadow model $B_{\text{cap}}$ (mitigating false positives from independent models). The only adjustment lies in the poisoning process for $P_{\text{cap}}$: Let $\mathcal{D}_{\text{cap}}$ denote the captioning dataset: $\mathcal{D}_{\text{cap}}=\{(\bm{x}_{i},\bm{c}_{i})\}_{i=1}^{N}$, where $\bm{c}_{i}\in\mathcal{C}$ is the ground-truth caption. $\gamma\%$ of samples with the lowest loss are selected as $\mathcal{D}_{s,\text{cap}}$, which is then poisoned by injecting a trigger $\bm{t}$ via generator $G(\cdot)$ to produce poisoned images $\bm{x}^{\prime}=G(\bm{x},\bm{t})$, with a target caption $\bm{c}_{t}$ ($e.g.$, “I don’t know.”) assigned to form $\mathcal{D}_{p,\text{cap}}$. $P_{\text{cap}}$ is obtained by fine-tuning $V_{\text{cap}}$ on $\mathcal{D}_{p,\text{cap}}\cup\mathcal{D}_{b,\text{cap}}$ ( where $\mathcal{D}_{b,\text{cap}}\triangleq\mathcal{D}_{\text{cap}}\setminus\mathcal{D}_{s,\text{cap}})$. The benign shadow model $B_{\text{cap}}$ follows the same construction as $B$ in Section 4.1.
Feature Projection and Output Difference Calculation. Since $V_{\text{cap}}$, $P_{\text{cap}}$, and $B_{\text{cap}}$ output text captions (not fixed-dimensional logits), we project these captions into a shared feature space using CLIP’s text encoder [radford2021learning]: $\mathcal{E}_{\text{CLIP}}:\mathcal{C}\to\mathbb{R}^{d}$, where $d$ is the embedding dimension. This encoder leverages pre-trained semantic alignment to ensure consistent comparison of text outputs. For $\bm{x}\in\mathcal{D}_{s,\text{cap}}$, let $\bm{c}_{V}=V_{\text{cap}}(\bm{x})$, $\bm{c}_{P}=P_{\text{cap}}(\bm{x})$, and $\bm{c}_{B}=B_{\text{cap}}(\bm{x})$ denote captions from the victim, poisoned, and benign models, respectively. These captions are projected into embeddings via $\mathcal{E}_{\text{CLIP}}$: $\bm{e}_{V}=\mathcal{E}_{\text{CLIP}}(\bm{c}_{V})$, $\bm{e}_{P}=\mathcal{E}_{\text{CLIP}}(\bm{c}_{P})$, and $\bm{e}_{B}=\mathcal{E}_{\text{CLIP}}(\bm{c}_{B})$. Output differences are computed as (mirroring Section 4.2): $\bm{O}_{v,\text{cap}}=\bm{e}_{V}-\bm{e}_{P}$, $\bm{O}_{b,\text{cap}}=\bm{e}_{B}-\bm{e}_{P}$, where $\bm{O}_{v,\text{cap}}$ captures $V_{\text{cap}}$’s dataset-specific features, and $\bm{O}_{b,\text{cap}}$ captures unrelated features.
The following process of training ownership meta-classifier and ownership verification follow Sections 4.2 and 4.3: the meta-classifier $C$ is trained on $\bm{O}_{v,\text{cap}}$ (label ‘+1’) and $\bm{O}_{b,\text{cap}}$ (label ‘-1’). For a suspicious model $S_{\text{cap}}$, project its caption $\bm{c}_{S}$ to $\bm{e}_{S}=\mathcal{E}_{\text{CLIP}}(\bm{c}_{S})$, compute $\bm{O}_{s,\text{cap}}=\bm{e}_{S}-\bm{e}_{P}$, and then apply the hypothesis test on $C(\bm{O}_{s,\text{cap}})$ to verify model ownership.
6.2.2 Experiment Settings
Dataset Selection. We conduct experiments on the COCO2014 dataset [lin2014microsoft], with a training set (16,000 image-caption pairs) and a validation set (8,000 image-caption pairs). For fine-tuning, hard distillation, and soft distillation attacks, 7,000 image-caption pairs are randomly selected from the validation set as the substitute dataset.
Settings of Personalization. The victim models are constructed by fine-tuning the pre-trained foundation models Qwen2.5-VL-7B-Instruct [^6] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct [bai2025qwen2] and Gemma-3-4B-IT [^7] https://huggingface.co/google/gemma-3-4b-it [team2025gemma] on the training set for 5 epochs. The prompt is “Describe this image”. Low-Rank Adaptation (LoRA) [hu2022lora] is adopted for efficient parameter tuning. Configurations of LoRA are provided in Appendix C.
Settings of Model Stealing. We simulate the model stealing attacks including: (1) Direct-copy: The adversary directly copies the parameters of the victim model; (2) Fine-tuning: The adversary obtains the stolen model by fine-tuning the victim model on the substitute dataset for 5 epochs; (3) Data Distill: We adopt the Align-KD method [feng2025align] (using its open-source implementation [^8] https://github.com/fqhank/Align-KD) and conduct the distillation attack using the COCO2014 training set. (4) Hard Distill: The stolen model is trained on the substitute dataset (with captions generated by the victim model) for 10 epochs. (5) Soft Distill: The stolen model is trained on the substitute dataset for 10 epochs, using both captions and visual token representations from the victim model, using the multimodal distillation strategy of LLaVA-KD [cai2025llava]. Regarding model architectures: for Direct-copy, Fine-tuning, Data Distill, and Soft Distill attacks, the stolen model adopts the same architecture as the victim model; for Hard Distill attacks, when Qwen2.5-VL-7B-Instruct serves as the victim model, the stolen model is Gemma-3-4B-IT, and when Gemma-3-4B-IT is the victim model, the stolen model is Qwen2.5-VL-3B-Instruct [bai2025qwen2]. In particular, we hereby exclude the Data-free Distill attack because, to the best of our knowledge, there exists no data-free knowledge distillation method tailored to image captioning, and directly extending existing techniques is ineffective. Current data-free distillation approaches are primarily developed for image classification [yu2023data, liu2024small, tran2024nayer], where class information naturally guides the synthesis of category-specific samples. In contrast, image captioning requires generating open-ended textual descriptions without categorical constraints, making these techniques unsuitable. Existing studies on vision–language distillation still depend on structured priors or paired data (e.g., image–caption pairs) to preserve semantic alignment [xuan2023distilling, wei2025open, cai2025llava]. Synthesizing image–caption pairs that align with distillation objectives in a data-free manner remains highly challenging.
Consistent with Section 5, we also include an “Independent” baseline. The independent baseline model is obtained by fine-tuning the foundation model on the COCO2014 training set for 10 epochs. For Qwen2.5-VL-7B-Instruct as the victim, the independent baseline is Gemma-3-4B-IT; for Gemma-3-4B-IT as the victim, the independent baseline is Qwen2.5-VL-7B-Instruct.
Settings of Defenses. For our method, detailed settings align with those in Section 5, with task-specific adaptations to image captioning. The poisoned shadow model $P_{\text{cap}}$ is constructed by fine-tuning the victim model $V_{\text{cap}}$ on $\mathcal{D}_{p,\text{cap}}\cup\mathcal{D}_{b,\text{cap}}$ for 5 epochs. We set the modification rate $\gamma\%=10\%$, resulting in $|\mathcal{D}_{s,\text{cap}}|=1,600$. We adopt a white square placed at the bottom-right corner of the image as the backdoor trigger, with the corresponding target caption set to “I don’t know”. The benign shadow model $B_{\text{cap}}$ is obtained by fine-tuning the foundation model for 5 epochs on $\mathcal{D}_{f,\text{cap}}$, filtered by dropping $\lambda\%=50\%$ of the lowest-loss samples from $V_{\text{cap}}$’s training set. The meta-classifier $C$ is trained for 100 epochs with a learning rate of 0.01.
Evaluation Metric. The hypothesis test setting follows Section 5.1. Evaluation uses confidence score $\Delta\mu$ and p-value, calculated via hypothesis testing on 100 sampled images ($m=100$). For stolen models, lower p-value and higher $\Delta\mu$ indicate stronger verification; for independent models, higher p-value and lower $\Delta\mu$ reflect better resistance to misclassification.
Table 14: Results on image captioning task.
[…]
$\Delta\mu$ p-value $\Delta\mu$ p-value
Direct-copy 0.90 $10^{-36}$ 0.89 $10^{-33}$
Fine-tuning 0.89 $10^{-20}$ 0.81 $10^{-31}$
Data Distill 0.89 $10^{-19}$ 0.86 $10^{-20}$
Hard Distill 0.80 $10^{-14}$ 0.79 $10^{-13}$
Soft Distill 0.75 $10^{-11}$ 0.78 $10^{-12}$
Independent -0.99 1.00 -0.98 1.00
6.2.3 Results
Table 14 presents the ownership verification results of Holmes on the COCO2014 image captioning task. For model stealing attacks, all strategies yield high $\Delta\mu$ (0.75–0.90) with p-values well below $\alpha$ ($10^{-11}$ to $10^{-36}$) across both victim models, indicating reliable ownership verification. For independent models, $\Delta\mu$ is negative (-0.98 to -0.99) and p-values are far above $\alpha$, confirming no false positives. These results validate Holmes’ effectiveness in image captioning, extending its practicality beyond classification tasks.
7 Potential Limitations and Future Directions
@@
Although our method demonstrates robust performance against various model stealing attacks (Section 5), we acknowledge several limitations that warrant discussion.
@@
Indirect Representation of Dataset-Specific Features. Our method utilizes the label-inconsistent backdoor attack to disrupt the dataset-specific features learned by the victim model. While our experiments show that this approach is effective, we acknowledge that other techniques could be explored for decoupling common features and disrupting dataset-specific ones.
@@
Computational Overhead. Our method involves fine-tuning a poisoned shadow model, a benign shadow model, and training an external meta-classifier, which introduces additional computational overhead. However, training shadow models for reference is a common practice in model watermarking and fingerprinting to evaluate whether these techniques impact model performance [jia2021entangled, maini2021dataset, wang2023free]. In our method, the fine-tuning cost of the shadow models is comparable to that of fine-tuning the victim model itself. For example, in the main experiments in Section 5, fine-tuning a shadow model on the CIFAR-10 dataset takes 1 hour and 58 minutes, while fine-tuning on the ImageNet dataset takes 1 hour and 7 minutes, both using an NVIDIA A100 GPU. Besides, this process can be accelerated through parallel GPU execution. The training cost for the meta-classifier is relatively mild, as its structure is lightweight and requires only a limited number of training samples. For example, training it on CIFAR-10 and ImageNet takes approximately 5 minutes on an NVIDIA A100 GPU.
@@
Generalization to Diverse Data Domains. While our method demonstrates effectiveness on standard vision benchmarks (CIFAR-10 and ImageNet), its robustness on specialized domains ($e.g.$, medical images or low-resolution images) remains unverified. The variations in data characteristics of these specialized domains, such as texture patterns or semantic granularity, might affect the decoupling of common and dataset-specific features, warranting further investigation.
8 Conclusion
@@
In this paper, we revisited the challenge of verifying ownership for personalized large vision models. We revealed that existing methods, which are typically designed for models trained from scratch, are either ineffective for fine-tuned models, introduce additional security risks, or are prone to misjudgment. To address these issues, we introduced a new fingerprinting paradigm and proposed a harmless model ownership verification method by decoupling common features. By creating a poisoned shadow model (retaining common features but disrupting dataset-specific ones via backdoor attacks) and a benign shadow model (learning distinct dataset-specific features), we distinguished the victim model’s dataset-specific features using output differences without altering the victim model or its training process. A meta-classifier and hypothesis test further enhanced verification reliability. We evaluated our method on the CIFAR-10 and ImageNet datasets, and the results demonstrated its effectiveness in detecting various types of model stealing attacks simultaneously.
Declarations
Data Availability. The experimental data that support the findings of this study are publicly available, including CIFAR-10 [krizhevsky2009learning], ImageNet [deng2009imagenet], and COCO2014 [lin2014microsoft].
Code Availability. Our codes are available at https://github.com/zlh-thu/Holmes.
Appendix A Proof of Theorem 1
We hereby provide the proof of our Theorem 1. In general, it is based on the property of $t$-distribution.
Definition 1.
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the posterior probabilities of the events $C(\bm{\bm{O}_{s}}(\bm{x}))=+1$ and $C(\bm{O}_{s}(\bm{x}))=-1$, respectively. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$-certainty) if and only if $H_{0}$ is rejected.
Theorem 1.
where $\Delta=t_{1-\alpha}^{4}+4t_{1-\alpha}^{2}(m-1)(\beta+\tau)(1-\beta-\tau)$, $t_{1-\alpha}$ is the $(1-\alpha)$-quantile of $t$-distribution with (m1)(m-1) degrees of freedom, and m is the sample size of $\bm{X}$.
Proof.
Since $\mu_{B}<\beta$, the original hypotheses $H_{0}:\mu_{B}+\tau=\mu_{S}$ and $H_{1}:\mu_{B}+\tau<\mu_{S}$ can be equivalently formulated as:
$H_{0}^{\prime}:\mu_{S}<\beta+\tau,$ (2)
$H_{1}^{\prime}:\mu_{S}>\beta+\tau.$ (3)
Let $E$ indicates the event of whether the suspect model $S$ can be identified by the meta-classifier $C$. As such,
$E\sim\mathcal{B}(1,p),$ (4)
where $p=\Pr(C(\bm{O}_{s}(\bm{X}))=+1)$ indicates $C$ identify $S$ as a stolen model and $\mathcal{B}$ is the Binomial distribution [hogg2005introduction].
Let $\bm{x_{1}},...,\bm{x_{m}}$ denote $m$ samples from $\mathcal{D}_{s}$ used for ownership verification and $E_{1},...,E_{m}$ denote their prediction events, we know that the identification success of $C$ rate $R$ satisfies
$R=\frac{1}{m}\sum_{i=1}^{m}E_{i},$ (5)
$R\sim\frac{1}{m}\mathcal{B}(m,p).$ (6)
According to the central limit theorem [hogg2005introduction], the identification success rate $a$ follows Gaussian distribution $\mathcal{N}(p,\frac{p(1-p)}{m})$ when $m$ is sufficiently large. Similarly, $(R-\beta-\tau)$ also satisfies Gaussian distribution. As such, we can construct the t-statistic as follows
$T\triangleq\frac{\sqrt{m}(R-\beta-\tau)}{s}\sim t(m-1),$ (7)
where $s$ is the standard deviation of $(R-\beta-\tau)$,
$s^{2}=\frac{1}{m-1}\sum_{i=1}^{m}(E_{i}-R)^{2}=\frac{1}{m-1}(m\cdot R-m\cdot R^{2}).$ (8)
To reject the hypothesis $H_{0}^{\prime}$ at the significance level $\alpha$, we need to ensure that
$\frac{\sqrt{m}(R-\beta-\tau)}{s}>t_{1-\alpha},$ (9)
where $t_{1-\alpha}$ is the $(1-\alpha)$-quantile of $t$-distribution with (m1)(m-1) degrees of freedom.
According to Equation 8 and Inequality 9, we have
$\sqrt{m-1}\cdot(R-\beta-\tau)-t_{1-\alpha}\cdot\sqrt{R-R^{2}}>0.$ (10)
To hold the Inequality 10, two conditions must be satisfied:
$R>\beta+\tau,$ (11)
and
$\sqrt{m-1}\cdot(R-\beta-\tau)>t_{1-\alpha}\cdot\sqrt{R-R^{2}}.$ (12)
The quadratic inequality of Inequality 12 is as:
$\small(m-1+t_{1-\alpha}^{2})R^{2}-(2(m-1)(\beta+\tau)+t_{1-\alpha}^{2})R+(m-1)(\beta+\tau)^{2}>0.$ (13)
For analyzing the valid interval of $R$, we consider the quadratic function as follow:
$\small f(R)=(m-1+t_{1-\alpha}^{2})R^{2}-(2(m-1)(\beta+\tau)+t_{1-\alpha}^{2})R+(m-1)(\beta+\tau)^{2}.$ (14)
The discriminant of $f(R)$ is given by $\Delta=t_{1-\alpha}^{4}+4t_{1-\alpha}^{2}(m-1)(\beta+\tau)(1-\beta-\tau)>0$. Thus, the two distinct real roots are given by:
$R_{1,2}=\frac{2(m-1)(\beta+\tau)+t_{1-\alpha}^{2}\pm\sqrt{\Delta}}{2(m-1+t_{1-\alpha}^{2})}.$ (15)
We can also find that $f(0)=(m-1)(\beta+\tau)^{2}>0$, $f(\beta+\tau)=t_{1-\alpha}^{2}(\beta+\tau)(\beta+\tau-1)<0$, and $f(1)=(m-1)(\beta+\tau-1)^{2}$.
According to the intermediate value theorem [andreescu2017intermediate], since $f(R)$ transitions from positive to negative in $(0,\beta+\tau)$, there must exist a root $R_{1}$ in this interval. Similarly, since $f(R)$ transitions from negative to positive in $(\beta+\tau,1)$, there must be a root $R_{2}$ in this interval. Thus, we have the strict order:
$0<R_{1}<(\beta+\tau)<R_{2}<1.$ (16)
Because $f(R)$ is positive for $R<R_{1}$ and $R>R_{2}$, and negative for $R_{1}<R<R_{2}$, it follows that the Inequality 13 is satisfied for $R<R_{1}$ or $R>R_{2}$. Given the additional constraint that $R>\beta+\tau$, the only valid solution is
$R>\frac{2(m-1)(\beta+\tau)+t_{1-\alpha}^{2}+\sqrt{\Delta}}{2(m-1+t_{1-\alpha}^{2})}.$ (17)
Appendix B Pseudocode of Holmes
The pseudocode of Holmes is in Algorithm 1.
Algorithm 1 The main pipeline of Holmes.
1:Input: Victim model $V$, foundation model $F$, suspicious model $S$, training set $\mathcal{D}$, shadow model modification rate $\gamma\%$, shadow model dropping rate $\lambda\%$, target class $y_{t}$, trigger pattern $\bm{t}$, poisoned image generator $G(\cdot)$, number of samples $m$ for the hypothesis test, and the significance level $\alpha$.
2:Output: Ownership verification result for suspicious model $S$.
3:Step 1: Create Shadow Models
4:(a) The Poisoned Shadow Model $P$:
5: Select $\gamma\%$ samples from $\mathcal{D}$ with the lowest loss values to create $\mathcal{D}_{s}$
6: Generate poisoned set: $\mathcal{D}_{p}=\{(\bm{x}^{\prime},y_{t})|\bm{x}^{\prime}=G(\bm{x};\bm{t}),(\bm{x},y)\in\mathcal{D}_{s}\}$
7: Fine-tune the victim model $V$ on $\mathcal{D}_{p}\cup\mathcal{D}_{b}$ to create the poisoned shadow model $P$, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$
8:(b) The Benign Shadow Model $B$:
9: Drop $\lambda\%$ samples from $\mathcal{D}$ with the lowest loss values to create $\mathcal{D}_{f}$
10: Fine-tune the foundation model $F$ on $\mathcal{D}_{f}$ to create the benign shadow model $B$
11:Step 2: Train Ownership Meta-Classifier $C_{w}$
12: Initialize the training set for meta-classifier $\mathcal{D}_{c}=\emptyset$
13:for each $(\bm{x},y)\in\mathcal{D}_{s}$ do
14:  Compute the output difference between $V$ and $P$: $\bm{O}_{v}(\bm{x})=V(\bm{x})-P(\bm{x})$
15:  Compute the output difference between $B$ and $P$: $\bm{O}_{b}(\bm{x})=B(\bm{x})-P(\bm{x})$
16:  $\mathcal{D}_{c}=\mathcal{D}_{c}\cup\{(\bm{O}_{v}(\bm{x}),+1)\}$
17:  $\mathcal{D}_{c}=\mathcal{D}_{c}\cup\{(\bm{O}_{b}(\bm{x}),-1)\}$
18:end for
19: Train the meta-classifier $C$ on the constructed set $\mathcal{D}_{c}$
20:Step 3: Ownership Verification
21: Randomly sample $m$ samples from $\mathcal{D}_{s}$
22: Compute the output difference between $S$ and $P$ for these $m$ samples: $\bm{O}_{s}(\bm{x})=S(\bm{x})-P(\bm{x})$
23: Conduct single-tailed pair-wise T-test using $\bm{O}_{s}(\bm{x})$ and $C$
24:if p-value $\alpha$ then
25:  Claim $S$ is stolen from $V$
26:else
27:  Claim $S$ is not stolen from $V$
28:end if
Appendix C LoRA Configurations of Experiments on Image Captioning
Detailed training configurations of LoRA are provided in Table 15.
Table 15: LoRA Fine-tuning Configuration
Parameter Name Value
Learning Rate $10^{-4}$
Warmup Ratio 0.05
Gradient Accumulation Steps 8
LoRA Rank 8
LoRA Alpha 32
Target Modules All Linear Layers
Freeze Vit True
Training Epochs 5